# ExpliCa: Evaluating Explicit Causal Reasoning in Large Language Models

### v3

@@ Appendix G Models’ Specifications
  • Qwen is a LLM family built by Alibaba Cloud. We used the $2.5$ version Yang et al. (2024) which underwent an optimization process during both the pre-training and post-training stages. This model is available at: https://huggingface.co/Qwen/Qwen2.5-7B-Instruct. For experiments on size effect, we used Qwen $2.5$ instruction-tuned models of increasing size, according to their number of parameters: $0.5$, $1.5$, $3$, $7$, $14$, and $32$ billions. [^26] The whole model family is available in this Hugging Face collection: https://huggingface.co/collections/Qwen/qwen25-66e81a666513e518adb90d9e.
  • Gemma is a model released by Google Team et al. (2024a) and it is built from the research and technology used to create Gemini models. We adopted version $2$ of this model Team et al. (2024b), with a size of $9$B parameters. This is the biggest among the small open models we adopted for our experiments. This model is available at: https://huggingface.co/google/gemma-2-9b-it.
@@
Appendix H Additional Models’ Results
@@
Appendix I Error Analysis
@@
Figure 13: Confusion matrices of the best performing models (GPT4o and Gemma) in prompting tasks, and the best performing model with APS (Falcon). The matrices show the errors by relation type (C. as Causality, T. as Temporal) and relation order (Ic. as Iconic, a.Ic. as Anti-Iconic).
Appendix J Correlation & Distribution
@@
Figure 14: Distribution of acceptability ratings for hu- mans and Gpt4o and Gemma (obtained with Outlines in zero-shot setting) and Falcon’s normalized perplexity.
Appendix K Size effect
Figure 15: The accuracy of Qwen models of different sizes on acceptability rating and APS is reported for each relation type and relation order, as well as overall.
Fig. 8 in Sec. 5.2 shows how Qwen models of different sizes perform according to two evaluation settings, accuracy on acceptability rating and APS. In Fig. 15 (Tab. 15 for the individual results), the overall accuracy on the acceptability rating task is divided into groups related to the relation condition of each item, according to the ground truth we obtained by human ratings. We observed that the same pattern as APS is also followed by temporal iconic relations, the most frequent condition in our dataset according to human ratings. temporal iconic relations are naturally the most frequent if we consider the chronological order of events in each narrative and the fact that a causal relation between events (most of the times) implies a temporal relation. Thus, the APS might rely more on the frequency of observation of a certain phenomenon in the dataset. On the contrary, the accuracy in the detection of relations in anti-iconic order seems not to follow a clear pattern.

### v4

@@ Appendix F Models Details
  • Qwen is a LLM family built by Alibaba Cloud. We used the $2.5$ version Yang et al. (2024) which underwent an optimization process during both the pre-training and post-training stages. This model is available at: https://huggingface.co/Qwen/Qwen2.5-7B-Instruct. For experiments on size effect, we used Qwen $2.5$ instruction-tuned models of increasing size, according to their number of parameters: $0.5$, $1.5$, 33, 77, $14$, and $32$ billions. [^26] The whole model family is available in this Hugging Face collection: https://huggingface.co/collections/Qwen/qwen25-66e81a666513e518adb90d9e.
  • Gemma is a model released by Google Team et al. (2024a) and it is built from the research and technology used to create Gemini models. We adopted version $2$ of this model Team et al. (2024b), with a size of 99B parameters. This is the biggest among the small open models we adopted for our experiments. This model is available at: https://huggingface.co/google/gemma-2-9b-it.
@@
Appendix G More Models’ Results
@@
Appendix H Error Analysis
@@
Figure 12: Comparison of different plots
Appendix I Correlation & Distribution
@@
Figure 13: This plot shows the
Appendix J Size effect
Figure 14: This plot shows the
@@
Fig. 7 in Sec. 5.3 shows how Qwen models’ of different sizes perform according to two evaluation settings, accuracy on acceptability rating and APS. In Fig. 14 (Tab. 14 for the individual results), the overall accuracy on the acceptability rating task is divided into groups related to the relation condition of each item, according to the ground truth we obtained by human ratings. We observed that the same pattern as APS is also followed by temporal iconic relations, the most frequent condition in our dataset according to human ratings. temporal iconic relations are naturally the most frequent if we consider the chronological order of events in each narrative and the fact that a causal relation between events (most of the times) implies a temporal relation. Thus, the APS might rely more on the frequency of observation of a certain phenomenon in the dataset. On the contrary, the accuracy in the detection of relations in anti-iconic order seems not to follow a clear pattern.